OpenStudy (anonymous):

need the domain and range for h(x)= - 1/2x-3

jimthompson5910 (jim_thompson5910):

domain: set of all real numbers (since you can plug in any number) range: set of all real numbers (since all linear functions can produce any number for a given x value)

jimthompson5910 (jim_thompson5910):

you can verify this with a graph

jimthompson5910 (jim_thompson5910):

this is assuming that the function is \[\large h(x)=-\frac{1}{2}x-3\]

OpenStudy (anonymous):

domain of this function is: 2x-3=0 x=3/2 cannot be D0m f(x) = all numbers except 3/2

OpenStudy (anonymous):

range of f(x) is all numbers

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

if the function is \[\large h(x)=\frac{1}{2x-3}\] then the domain is \[\large \{x|x\in\mathbb{R},x\neq\frac{3}{2}\}\] and the range is \[\large \{y|y\in\mathbb{R},y\neq0\}\]

OpenStudy (anonymous):

yes the function is the one you just posted but with a negative in front of the fraction

jimthompson5910 (jim_thompson5910):

oh right, my bad, but the domain and range are still the same

OpenStudy (anonymous):

thank you